The following statement uses a numbering system that is NOT decimal: 235 + 32 = 311....

  1. The following statement uses a numbering system that is NOT decimal: 235 + 32 = 311. What “base” does this numbering system have? Hint, try using trial and error.
  1. Calculate the equation: 2 - 5 = -3 using 4-bit signed 2's complement math. Show all your work. Prove the answer -3 using 2’s complement.
  1. What is the lowest number and the highest number that can be represented with 5 bits using:
    1. unsigned numbers?
    2. 2's complement numbers?
Answer #1

235 + 32 =311

here base is 5

the value of 235 base 5 to decimal is 95 and 32 is 20 and 95 + 20=115 in decimal and 211 in base 5

2-5=-3

2= 0010

5= 0101

-5=1011

2-5=1101

since ist bit is 1 so it is negative

2's complement of 1101=0011=3

ans is -3

highest value and lowest

unsigned :- 11111=31

00000=0

in 2's

highest value 01111=15

lowest 11111=-15